# Toward Quantum-Aware Machine Learning: Improved Prediction of Quantum Dissipative Dynamics via Complex Valued Neural Networks

### v1

@@ IV Results and Discussion
For each system, we independently trained two neural network models—a RVNN and a CVNN—using identical training data. Both models adopt a fully connected sequence-to-sequence architecture consisting of five layers: one input layer, three hidden layers, and one output layer. In the RVNN, all layers employ real-valued weights and biases together with the ReLU activation function. In contrast, the CVNN is constructed using custom complex-valued linear layers with complex weights and biases, combined with a complex ReLU (CReLU) activation function that applies the ReLU operation separately to the real and imaginary components. To ensure a fair comparison, the total number of trainable parameters in the CVNN and RVNN architectures is kept approximately the same. Moreover, both network architectures are implemented from scratch within the PyTorch framework, allowing full control over the operations. A schematic overview of the training and optimization workflow for both models is shown in Fig. 1.
@@
For the 7-site and 8-site FMO complexes, training data are likewise sourced from the QD3SET-1 database. These datasets describe excitation energy transfer dynamics initiated at site 1 and site 6 for the 7-site complex, with site 8 additionally included for the 8-site system. The associated parameter space $(\lambda,\gamma,T)$ spans bath reorganization energy, relaxation rate, and temperature, yielding a total of 500 simulations for each initial excitation scenario. The quantum dynamics are propagated using a trace-conserving local thermalizing Lindblad master equation (LTLME)[mohseni2008environment]. The system Hamiltonians are parameterized following the models of Adolphs and Renger for the 7-site FMO complex[adolphs2006proteins] and Jia et al. for the 8-site extension[jia2015hybrid].
@@
To improve training efficiency and ensure representative coverage of the parameter space, farthest point sampling[dral2019mlatom, ullah2022predicting] is employed to select a subset of trajectories from each dataset. For the SB model, we consider only the symmetric case, comprising 500 trajectories from $\mathcal{D}_{\mathrm{sb}}$. For each of the four systems—the SB model and the three FMO prototypes—we then select 400 trajectories for training via farthest point sampling, with the remaining trajectories reserved for out-of-sample testing.
@@
4-site FMO complex 6.30e-05 7.21e-04
7-site FMO complex 4.02e-05 6.36e-05
8-site FMO complex 4.98e-05 5.33e-05
@@
We now turn to the physical fidelity of the predicted RDMs. We first assess trace conservation by comparing the CVNN and RVNN predictions. Figure 3 shows the trace evolution for a representative test trajectory, while Table 1 reports the mean absolute error (MAE) of the trace relative to unity, averaged over 100 trajectories.
@@
4-site FMO complex 39.46 43.23
7-site FMO complex 40.33 71.36
8-site FMO complex 57.43 62.69
@@
For the low-dimensional SB model, both CVNN and RVNN exhibit nearly identical trace behavior in the single-trajectory example and comparable ensemble-averaged MAEs on the order of $10^{-5}$. As shown in Table 1, clear differences emerge for the FMO systems. In the 4-site FMO complex, the CVNN achieves substantially lower trace error than the RVNN, with the latter displaying an order-of-magnitude larger deviation. For the 7-site and 8-site FMO complexes, both models maintain trace errors in the $10^{-5}$ range; however, the CVNN consistently yields slightly smaller MAEs
@@
For the SB model, both CVNN and RVNN preserve positive semi-definiteness in all predictions, with no negative eigenvalues observed in either the single-trajectory spectra or the ensemble-averaged statistics. A similar level of consistency is found for the 4-site FMO complex, where both models exhibit only a small number of negative eigenvalues on average, with the CVNN showing a modest but systematic reduction relative to the RVNN.
@@
SB model 6.03e-4 (1.91e-3, 5.26e-4) 7.85e-4 (1.95e-3, 6.55e-4)
4-site FMO 1.52e-3 (9.16e-4, 7.60e-5) 2.75e-3 (1.80e-3, 2.12e-4)
7-site FMO 6.29e-4 (5.28e-4, 6.56e-5) 1.23e-3 (1.01e-3, 9.24e-5)
8-site FMO 5.13e-4 (6.98e-4, 6.07e-5) 5.78e-4 (8.08e-4, 7.82e-5)
As the system size increases, however, more pronounced differences emerge. For the 7-site and 8-site FMO complexes, violations of positive semi-definiteness become more frequent, reflecting the increased complexity and coherence of the dynamics. For these systems, CVNN consistently produces fewer negative eigenvalues than the RVNN, both in the representative eigenvalue spectra shown in Fig. 4 and in the ensemble-averaged results summarized in Table 2. In particular, the average number of negative eigenvalues is reduced by nearly a factor of two for the 7-site system and remains systematically lower for the 8-site complex.
@@
To further evaluate the CVNN and RVNN models, we compare their predictive accuracy for RDM dynamics. Figure 5 shows the time evolution of diagonal (population) and off-diagonal (coherence) elements for a representative test trajectory. For clarity, both diagonal and off-diagonal elements are shown for the SB model, while for the FMO complexes the off-diagonal dynamics are provided in the Supporting Information. Ensemble-averaged accuracy is quantified in Table 3, which reports the MAE averaged over 100 independent test trajectories and all time steps, separately for diagonal and off-diagonal RDM elements.
For the SB model, both architectures achieve comparable accuracy for populations and coherences, consistent with the relatively simple dynamics and the limited role of long-lived quantum coherence. As system size increases, systematic differences emerge. For the 4-site and 7-site FMO complexes, the CVNN consistently yields lower MAEs than the RVNN for both diagonal and off-diagonal elements, with particularly pronounced improvements in the coherences. In the 7-site FMO case, this enhanced accuracy correlates with the CVNN’s superior preservation of positive semi-definiteness, as evidenced by the reduced number of negative eigenvalues reported in Table 2.
For the 8-site FMO complex, both models achieve similar accuracy for the population dynamics, while the CVNN maintains a modest but systematic advantage in predicting off-diagonal elements. Notably, despite comparable MAEs, the RVNN exhibits larger violations of trace conservation and positive semi-definiteness for this system, highlighting that numerical accuracy alone does not guarantee physical fidelity.
Overall, these results demonstrate that CVNNs not only match or exceed RVNNs in predictive accuracy but also offer improved robustness with respect to fundamental quantum constraints, particularly as system size and coherence complexity increase.
@@ V Concluding Remarks
VI Acknowledgments
A.U. acknowledges funding from the National Natural Science Foundation of China (No. W2433037) and the Natural Science Foundation of Anhui Province (No. 2408085QA002).
VII Data availability

### v2

@@ IV Results and Discussion
For each system, we independently trained two neural network models—a RVNN and a CVNN—using identical training data. Both models adopt a fully connected sequence-to-sequence architecture consisting of five layers: one input layer, three hidden layers, and one output layer. In the RVNN, all layers employ real-valued weights and biases together with the ReLU activation function. In contrast, the CVNN is constructed using custom complex-valued linear layers with complex weights and biases, combined with a complex ReLU (CReLU) activation function that applies the ReLU operation separately to the real and imaginary components. To ensure a fair comparison, we capped the number of trainable parameters in the CVNN at a level approximately equal to or below that of the RVNN. Detailed architectures are listed in Tables S1 and S2 (Section S2 of the Supporting Information). Moreover, both network architectures are implemented from scratch within the PyTorch framework, allowing full control over the operations. A schematic overview of the training and optimization workflow for both models is shown in Fig. 1.
@@
For the 7-site and 8-site FMO complexes, training data are likewise sourced from the QD3SET-1 database. These datasets describe excitation energy transfer dynamics initiated at site 1 and site 6 for the 7-site complex, with site 8 additionally included for the 8-site system. The associated parameter space $(\lambda,\gamma,T)$ spans bath reorganization energy, relaxation rate, and temperature, yielding a total of 500 simulations for each initial excitation scenario. The quantum dynamics are propagated using a trace-conserving local thermalizing Lindblad master equation (LTLME)[mohseni2008environment]. The system Hamiltonians are parameterized following the models of Adolphs and Renger for the 7-site FMO complex[adolphs2006proteins] and Jia et al. for the 8-site extension[jia2015hybrid]. A detailed discussion of the parameter ranges and sampling scheme is provided in Section S2 of the Supporting Information.
@@
To improve training efficiency and ensure representative coverage of the parameter space, farthest point sampling[dral2019mlatom, ullah2022predicting] is employed to select a subset of trajectories from each dataset (See Section S2 of the Supporting Information). For the SB model, we consider only the symmetric case, comprising 500 trajectories from $\mathcal{D}_{\mathrm{sb}}$. For each of the four systems—the SB model and the three FMO prototypes—we then select 400 trajectories for training via farthest point sampling, with the remaining trajectories reserved for out-of-sample testing.
@@
4-site FMO complex 1.06e-04 2.09e-04
7-site FMO complex 5.48e-05 8.73e-05
8-site FMO complex 2.23e-05 8.52e-05
@@
We now turn to the physical fidelity of the predicted RDMs. We first assess trace conservation by comparing the CVNN and RVNN predictions. Fig. 3 shows the trace evolution for a representative test trajectory, while Table 1 reports the mean absolute error (MAE) of the trace relative to unity, averaged over 100 trajectories.
@@
4-site FMO complex 38.05 40.63
7-site FMO complex 60.17 67.85
8-site FMO complex 53.87 73.71
@@
For the low-dimensional SB model, both CVNN and RVNN exhibit comparable trace MAEs on the order of $10^{-5}$. As shown in Table 1, the CVNN demonstrates a clear advantage for the FMO complexes, which becomes more systematic as the system representation becomes more complete. The 4-site FMO complex (comprising the four prominent sites of the full 7-site system) already shows a twofold trace accuracy improvement for the CVNN over the RVNN. For the complete 7-site complex, the CVNN maintains a 1.6x lower error. Most notably, for the 8-site complex, the CVNN’s error reduces to $2.23\times 10^{-5}$, which is nearly 4 times smaller than the RVNN’s error ($8.52\times 10^{-5}$).
@@
For the SB model, both CVNN and RVNN preserve positive semi-definiteness, with zero negative eigenvalues in the ensemble average. In contrast, violations emerge for the FMO complexes. For the 4-site FMO complex, both models exhibit a similar scale of violations, with the CVNN showing a reduction of approximately 6% relative to the RVNN baseline (38.05 vs. 40.63). The advantage of the CVNN becomes more pronounced with increasing system complexity. For the 7-site FMO complex, the CVNN produces  11% fewer negative eigenvalues than the RVNN (60.17 vs. 67.85). This performance gap widens substantially for the 8-site complex, where the CVNN achieves a reduction of approximately 27% relative to the RVNN (53.87 vs. 73.71).
@@
SB model 7.01e-4 (2.28e-3, 6.51e-4) 7.85e-4 (1.17e-3, 6.55e-4)
4-site FMO 1.94e-3 (1.17e-3, 7.62e-5) 2.87e-3 (1.78e-3, 9.44e-5)
7-site FMO 5.23e-4 (4.64e-4, 6.55e-5) 1.23e-3 (1.07e-3, 8.73e-5)
8-site FMO 4.87e-4 (6.53e-4, 6.05e-5) 1.40e-3 (1.68e-3, 9.36e-5)
@@
To further evaluate the CVNN and RVNN models, we compare their predictive accuracy for RDM dynamics. Fig. 5 shows the time evolution of diagonal (population) and off-diagonal (coherence) elements for a representative test trajectory. For clarity, both diagonal and off-diagonal elements are shown for the SB model, while for the FMO complexes the off-diagonal dynamics are provided in the Supporting Information (Section S3). Ensemble-averaged accuracy is quantified in Table 3, which reports the MAE averaged over 100 independent test trajectories and all time steps, separately for diagonal and off-diagonal RDM elements. A time resolved MAE trend is shown in Figs. S2 and S3 (Section S4 of the Supporting Information)
For the SB model, predictive accuracy is comparable between architectures. The CVNN reduces diagonal MAE by approximately 11% (relative to the RVNN baseline), while the RVNN yields a  49% lower MAE for the real coherence components (relative to the CVNN). These offsetting advantages indicate that neither model holds a decisive edge for this minimally coherent system. However, as system complexity increases, a clear CVNN advantage emerges. For the 4‑site FMO complex, the CVNN shows a consistent reduction in MAE: approximately 32% lower for diagonal elements and 34% lower for the real parts of coherences compared to the RVNN (relative to the RVNN baseline).
This advantage becomes more pronounced for the larger systems. In the 7‑site FMO complex, the CVNN’s diagonal error is less than half that of the RVNN (5.23e‑4 vs. 1.23e‑3), and its error for the real part of coherences is  57% lower (4.64e‑4 vs. 1.07e‑3). The trend continues for the 8‑site FMO complex, where the CVNN again yields substantially lower errors: roughly 65% lower for diagonals and 61% lower for the real coherence components. Together, these results demonstrate that the CVNN’s accuracy advantage—especially in capturing populations and key coherences—grows systematically as the open quantum system increases in size and coherence.
To demonstrate the general applicability and superior performance of the CVNN framework, we extend our analysis to the 8-site FMO complex modeled with an Ohmic spectral density. The results, presented in Supplementary Information (Section S5), corroborate our primary findings and confirm the consistent advantage of CVNNs over RVNNs across different environmental models.
We further verified the robustness of the CVNN’s superior performance by testing three different random seeds (42, 100, and 123). As detailed in Section S6 of the Supporting Information, the CVNN outperformed the RVNN across all metrics—training/validation loss, trace conservation, average negative eigenvalues, and RDM prediction accuracy—for every seed tested. This seed-invariance analysis confirms that the CVNN’s enhanced accuracy and physical consistency are not artifacts of initialization but are intrinsic to the complex-valued architecture.
Overall, all these results demonstrate that CVNNs not only match or exceed RVNNs in predictive accuracy but also offer improved robustness with respect to fundamental quantum constraints, particularly as system size and coherence complexity increase.
@@ V Concluding Remarks
VI SUPPLEMENTARY MATERIAL
The supplementary material, provided as a separate file, includes additional results and further details on model architectures, parameters range and the farthest-point sampling strategy, the dynamics of off-diagonal elements (corresponding to Fig. 5), a time-resolved MAE analysis, a robustness test on 8-site FMO complex with Ohmic spectral density, and a seed-invariance study.
VII Acknowledgments
A.U. acknowledges funding from the National Natural Science Foundation of China (No. W2433037) and the Natural Science Foundation of Anhui Province (No. 2408085QA002). M.Y. acknowledges funding from Scientific Research Projects of Anhui Provincial Department of Education under Grant Nos. 2025AHGXZK20126 and 2025AHGXZK50054.
VIII Data availability